Weewee Inc, will make 10,000 special edition weewee chairs. Each has a direct variable cost of $88/unit, + sales commissions of 5% and import taxes of 7% the sales price. The fixed costs were $300,000. What is the minimum selling price?
Answer #1

Total sales price = direct variable cost + fixed costs + sales commissions + import taxes

Total sales price = 88*10000 + 300000 +Total sales price*12%

0.88 Total sales price = 1180000

Total minimum sales price = 1340909

minimum selling price = 1340909 / 10000 = $134 per unit